Consider two Tanzanian Banks, A and B, with the following quotes which are expressed as Ksh/Tsh.

Bank A Bank B
Tsh quote Bid Ask Bid Ask
Ksh0.06 Ksh.0.07 Ksh 0.08 Ksh0.09
Required
Compute the gain from locational arbitrage to a Tanzanian investor with Kshs.30 million to invest (5 marks)
b) A 90 days treasury bill with a par value of Kshs 200 and auctioned at Kshs 194. Find the discount rate and then investment yield or rate. (2 marks
a) If the price level as measured by the CPI was 150 in 2018 and the project price level for 2019 is 120. Calculate the expected inflation rate between 2018 and 2019. (1 marks)
1
Expert's answer
2019-11-29T10:03:46-0500

a) The gain is:

Gain = (Bid B - Ask A)×Investment = (0.08 - 0.07)×5,000,000 = Kshs 500,000.

b) Yd = (200 - 194)/200×(360/90)×100% = 12%.

The investment yield is: 0.12×200/194 = 0.1237 or 12.37%.

c) The expected inflation rate between 2018 and 2019 is: (150 - 120)/120×100% = 25%.